Direct Materials Mix and Yield Variances

Chypre, Inc., produces a cologne mist using a solvent mix (water and pure alcohol) and aromatic compounds (the scent base) that it sells to other companies for bottling and sale to consumers. Chypre developed the following standard cost sheet:

Direct Material Mix Mix Proportion SP Standard Cost
Solvent mix 855 gallons 0.95     $5.55 $4745.25
Aromatic compounds 45 0.05 8040.00 361800.00
   Total 900 gallons $366,545.25
Yield 837 gallons

On May 2, Chypre produced a batch of 1,200 gallons with the following actual results:

Direct Material Actual Mix
Solvent mix 1,135.2 gallons
Aromatic compounds 64.8
   Total 1,200 gallons
Yield 1,080 gallons

Required:

Instructions for parts 3 and 4: If a variance is zero, enter "0" and select "Not applicable" from the drop down box.

1. Calculate the yield ratio. If required, round your answer to two decimal places.

2. Calculate the standard cost per unit of the yield. If required, round your answer to the nearest cent.
$ per unit of the yield

3. Calculate the direct materials yield variance. If required, round your answer to the nearest cent.
$  Unfavorable

4. Calculate the direct materials mix variance. If required, round your answer to the nearest cent.
$  Unfavorable
Answer #1

Solution:

1. Yield Ratio= Standard Qty/Standard Yield*Actual Yield

Solvent Mix=855/837*1080 = 1103.23

Aromatic Compounds= 45/837*1080 =58.06

2. Standard cost per unit of yield

Solvent Mix= 1103.23* 5.55 =6122.93

Aromatic Compounds =58.06*8040 = 466802.40

3. Direct Material Mix Variance= (Revised Std Qty-Actual Qty)*Standard Price

Solvent Mix=(1103.23-1135.20)* 5.55=177.49 Unfavorable

Aromatic Compounds= (58.06-64.8)* 8040=54186.4 Unfavorable

4. Direct Material Yield Variance = (Standard Qty- Revised Standard Qty)*Standard Price

Solvent Mix=(855-1103.23)* 5.55 = 1377.68 Unfavorable

Aromatic Compounds= (45-58.06)* 8040 = 105002.40 Unfavorable